HETAV PATEL

Grade 11,

Q)​A particle moves along the x-direction with constant acceleration.The displacement,measured from a convenient position,is 2m at time t=0 and is zero when t=10s.If the velocity of the particle is momentary zero when t=6s,determine the acceleration a and the velocity v when t=10s. PLZ SOLVE IT SIR.......................

Q)​A particle moves along the x-direction with constant acceleration.The displacement,measured from a convenient position,is 2m at time t=0 and is zero when t=10s.If the velocity of the particle is momentary zero when t=6s,determine the acceleration a and the velocity v when t=10s.
PLZ SOLVE IT SIR....................... 

Grade:11

2 Answers

atreyee
9 years ago
Let the initial velocity be u and acceleration be a. The displacement of the particle from t=0 to t=10s is -2m.
hence, 10u+50a=-2 and u+6a=0 (using the formulae ut+1/2at2 = s and u+at=v)
Solving these two equations, we get a=0.2m/s2 and u=-1.2m/s
at t=10s, the velocity v=-1.2+10*0.2=0.8m/s
Bandi Gayatri
6 years ago
Since the x-t graph is parabolic. Write the equation of a parabola. With vertex as (6,y). Find the equation of the parabola by substituting the values given. That is (10,0) & (0,2) . on getting the equation of the parabola, differentiate once and twice for velocity and acceleration respectively.